OpenStudy (anonymous):

Which group of points are all on the line y = -2x + 1?

OpenStudy (anonymous):

A (3, –5); (1, –1); (2, –3) B(3, –7); (1, –3); ( 2, –3) C (3, –7); (1, –1); ( 2, –5) D(3, –5); (1, –3); (2, –5)

jimthompson5910 (jim_thompson5910):

Any ideas where to start?

OpenStudy (anonymous):

nopee

jimthompson5910 (jim_thompson5910):

Let's say that choice C was the answer (this is hypothetical for now since we don't know yet) If it was, then (3, –7), (1, –1), and ( 2, –5) would be solutions If (3,-7) was a solution, then x = 3 and y = -7, so... y = -2x + 1 -7 = -2(3) + 1 -7 = -6 + 1 -7 = -5 ... which is false. So (3, –7) is NOT a solution, which means choice C is NOT the answer. Do the same with the other answer choices to see if those points are solutions.